# Gradient-flow adaptive importance sampling for Bayesian leave one out cross-validation for sigmoidal classification models

### v1

@@ 5 Experiments
5.1 Implementation
We implemented our methodology in Python using the github:mederrata/bayesianquilts wrapper for Tensorflow probability [Dillon et al., 2017]. For Pareto smoothing, we used the Python implementation available at github:avehtari/PSIS.
5.2 Dataset and model
@@
For demonstration, we used a public domain ovarian cancer micro-array dataset Hernández-Lobato et al. [2010], Schummer et al. [1999]. This dataset consists of $n=54$ observations of $p=1056+1$ predictors. As an example of a $p\gg n$ problem, model-agnostic $1/\ell$ importance sampling is insufficient for computing LOO expectations. Notably, Paananen et al. [2021] used this dataset to test their moment-matching adaptive importance sample where they successfully decreased the number of observations where $\hat{k}>0.7$ from between 34 and 36 to between 11 and 20, depending on the number of posterior samples. We reproduced their logistic regression model, using the same regularized-horseshoe [Piironen and Vehtari, 2017b, c, Carvalho et al., 2009] prior, and the same statistical inference scheme using Stan, which we interfaced to Python using the package cmdstanpy. We ran four parallel Markov Chains, with twelve thousand burn-in iterations, retaining 2000 samples per chain after thinning to every other sample.
Prior to any adaptation, we estimated $\hat{k}$ for the corresponding $1/\ell$ tail weights for this model. Repeatedly simulating 64 posterior draws, we found $27\pm 4$ observations having $\hat{k}$ in excess of the stability cutoff.
5.3 Decreasing $\hat{k}$
We scanned different values of $\rho=10^{-r}$, for $r\in\{0,1,\ldots,6\}$, evaluating all three transformations for a given value of $\rho.$ We found that for $21\pm 4$ of the initially $\hat{k}>0.7$ observations that at least one of the transformations reduced from super-threshold to sub-threshold in $\hat{k}.$ Fig. 2 shows the results of one such simulated posterior draw. Generally, for too-large a value of $\rho,$ all $\hat{k}$ values increased.
Figure 2: Applying transformations (blue=KL descent, green=variance descent, purple=LL descent) for different values of $\rho$ (point labels are $-\log_{10}\rho$) to posterior samples of a logistic regression model fitted to the ovarian cancer dataset and evaluating $\hat{k}$ for those observations where the untransformed 1/$1/\ell$ IS weights (Eq. 7) have $\hat{k}>0.7$.
5.4 Area under ROC/PRC
Since $n\ll p,$ logistic regression is easily able to perfectly fit the dataset as defined by achieving in-training AUROC/AUPRC scores of approximately one.
For the model trained using MCMC, the LOO AUROC and AUPRC curves have areas slightly less than one (Fig. 2(a)). In contrast, the model, when inferred using mean-field ADVI, has lower LOO AUROC and AUPRC scores. This finding is likely due to the fact that the MCMC model, when using the full posterior, captures associations between the model parameters – in effect finding a lower-dimensional “factorization” of the features. The mean-field model is unable to capture parameter correlations by design and its LOO performance suffers.
(a) MCMC model
(b) Mean-field ADVI model
@@ 6 Discussion
In this manuscript we introduced an adaptive importance sampler for using pre-trained full-data posteriors to approximate leave one out cross validation (LOO) in Bayesian classification models. The objective of adaptation to bring the sampling distribution (the full data posterior) closer to the target LOO posterior distributions for each data point. The methodology is based on taking a single step according to the gradient flow corresponding to minimization of a given objective. We introduce two such transformations: KL divergence descent and variance descent, along with a simple log-likelihood descent transformation. We presented explicit formulae for these transformations for logistic regression and ReLU-activated artificial neural networks with one hidden layer – the latter by computing the exact spectral decomposition of its Hessian matrix. We described how one can easily approximate the Jacobian of the transformations for more-complicated models, including for ReLU neural networks of any size. The adaptive importance sampler is ultimately used to estimate the expected LOO prediction for each given datapoint – quantities that can be used to compute downstream model generalization metrics such as ROC/PRC curves and the area under these curves.
6.1 Limitations
@@
The main tradeoff of this method versus the model-agnostic PSIS-LOO method is that this method is model-dependent. In order to use this methodology for a given model, one needs to be able to evaluate gradients of the model with respect to parameters – and also the gradients of the corresponding prior distribution. Both the KL descent and variance descent transformations require computing the the posterior density – when a variational approximation of the posterior is not available or trustworthy this computation is costly for large datasets.
6.2 Extensions
@@
In this manuscript we focused on classification problems but the methodology for adapting the importance sampler is much broader. In the Supplemental Materials one may find more-general formulae for the KL and variance descending transformations. In medical and industrial contexts, one is often interested in whether an individual or unit will experience an outcome within a certain time interval. For instance, policymakers are interested in hospital readmission within 30 days post discharge [Xia et al., 2023, Chang et al., 2023] because these readmissions are possibly preventable. In these problems, one may apply survival modeling to characterize the lifetime distribution, and additionally evaluate a model according to its classification performance at a given cut-off time $T$. Our methodology can easily be used for assessing such models.
@@
Another extension to this methodology is to take more steps along the gradient flow for a given objective. Finally, we resolved the spectrum of the Hessian matrix for shallow ReLU models – the spectral decomposition for larger ReLU models may be useful for other analyses.
{contributions}
JCC conceived the idea, developed the theory, wrote the code, and drafted the paper XL, SX, CC developed the theory and edited the paper. HRY and JP edited the paper.
Acknowledgements.

### v2

@@ 5 Experiments
PDF versions of Jupyter notebooks for producing the results in the text are included in the Supplemental Results. As baselines for comparison, we also evaluated generalizations of the moment matching (MM1/MM2) affine transformation method of Paananen et al. [2021] where we used a pre-factor of $\rho$ to control the magnitude of the change in the transformation (details in  S.3), and the log-likelihood (LL) gradient descent method of Elvira et al. [2022] (derivations of this transformation for sigmoidal models are available in  S.1).
@@
Dataset and model: For demonstration, we used a public domain ovarian cancer micro-array dataset Hernández-Lobato et al. [2010], Schummer et al. [1999]. This dataset consists of $n=54$ observations of $p=1056+1$ predictors. As an example of a $p\gg n$ problem, model-agnostic $1/\ell$ importance sampling is insufficient for computing LOO expectations. Notably, Paananen et al. [2021] used this dataset to test their moment-matching adaptive importance sampler where they successfully decreased the number of observations where $\hat{k}>0.7$ from approximately $35$ to approximately $20$ using $s=1000$ posterior samples. We reproduced their logistic regression model, using the same regularized-horseshoe [Piironen and Vehtari, 2017b, c, Carvalho et al., 2009] prior, and the same statistical inference scheme within Stan, which we interfaced to Python using the package cmdstanpy. We ran four parallel Markov Chains, with twelve thousand burn-in iterations, retaining 2000 samples per chain (more details available in  S.4.1). We then evaluated the transformation methods on resamplings of the retained MCMC samples.
Adaptation:
KL Var LL MM1 MM2 #u̇nsuccessful
$31.5\pm 2.8$
$16.2\pm 3.0$
$18.2\pm 2.6$
$12.41\pm 2.6$
$20.4\pm 3.1$
$2.9\pm 1.6$
$3.1\pm 3.1$
$0.8\pm 1.0$
$0.2\pm 0.1$
Table 1: Counts of unsuccessful adaptations (mean $\pm$ standard deviation) when using at least one of the given combination of transformations across the step sizes $\rho\in\{4^{-r}:r\in\{0,1,\ldots,10\}\}$, as seen in one hundred simulations of parameter sample size $s=1000.$ Lower is better.
[…]
We scanned different values of $\rho=4^{-r}$, for $r\in\{0,1,2,\ldots,10\}$, evaluating both transformations (KL/Var), and the comparison methods (LL/MM1/MM2) for a given value of $\rho.$ We performed this procedure $100$ times, using samples of size $s=1000.$ Recall that adaptation is successful if any of the considered transformations can reduce $\hat{k}$ to below $0.7$.
Table 1 presents statistics (mean $\pm$ standard deviation) for the number of observations where adaptation fails when using the given combination of methods. For a representative instance of the simulation procedure, Fig. 2 depicts the computed values of $\hat{k}$ obtained, organized by the index of each relevant observation within the dataset. The large black dot is the original $\hat{k}$ value and the smaller colored dots are the post-transformation KL and Var values for varying values of $\rho$. Fig. 3 shows the corresponding LOO- receiver operator characteristic (ROC) and precision recall curve (PRC) obtained by using the transformation that resulted in the best $\hat{k}$ value for each observation, predicting the LOO estimate of predictive probability, and feeding those probabilities into the relevant formulae for computing ROC and PRC. We provide these curves for both MCMC and ADVI -inferred variants of the model.
@@ 6 Discussion
In this manuscript, we introduced an adaptive importance sampler for using pre-trained full-data posteriors to approximate leave one out cross validation (LOO CV) in Bayesian classification models. The objective of importance-sampling LOO (IS LOO) is to compute LOO CV without incurring the computational cost of refitting a given model. The objective of adaptation to bring the sampling distribution (the full data posterior) closer to the target LOO posterior distributions for each data point so that IS LOO produces reliable estimates.
@@
Our methodology is based on taking a single step according to the gradient flow corresponding to the minimization of a given objective. We introduce two such transformations: KL divergence descent and variance descent. We presented explicit formulae for these transformations for logistic regression and ReLU-activated artificial neural networks with one hidden layer – the latter by computing the exact spectral decomposition of its Hessian matrix. An analysis of the same dataset using a Bayesian ReLU-net is available in S.4.2. We described how one can easily approximate the Jacobian of the transformations for more-complicated models, including for ReLU neural networks of any size. The adaptive importance sampler is ultimately used to estimate the expected LOO prediction for each given datapoint – quantities that can be used to compute downstream model generalization metrics such as ROC/PRC curves and the area under these curves.
Contrasting and synergizing methods: Examining Tabel 1, taken individually, the KL and Var gradient flow-based transformations perform comparably to the original MM transformations evaluated in Paananen et al. [2021]. However, the generalized MM1/MM2 transformations have by-far the best performance in shifting $\hat{k}$. Yet these two MM methods, used either alone or together, usually do not completely get the job done. Each unsuccessful adaptation means that the model must be refit one additional time at high computational cost. Fortunately, using all the evaluated methods in-unison resulted in successful adaptation for all data points most of the time. The general strategy is then to loop through observations and try successive transformations for each observation until adaptation is successful.
@@
Limitations: The main tradeoff of this method versus the model-agnostic PSIS-LOO method is that this method is model-dependent. In order to use this methodology for a given model, one needs to be able to evaluate gradients of the model with respect to parameters – and also the gradients of the corresponding prior distribution. Both the KL descent and variance descent transformations require computing the the posterior density – when a variational approximation of the posterior is not available or trustworthy this computation is costly for large datasets.
@@
Extensions: In this manuscript, we focused on classification problems but the methodology for adapting the importance sampler is much broader. In the Supplemental Materials one may find more-general formulae for the KL and variance descending transformations. In medical and industrial contexts, one is often interested in whether an individual or unit will experience an outcome within a certain time interval. For instance, policymakers are interested in hospital readmission within 30 days post discharge [Xia et al., 2023, Chang et al., 2023] because these readmissions are possibly preventable. In these problems, one may apply survival modeling to characterize the lifetime distribution, and additionally evaluate a model according to its classification performance at a given cut-off time $T$. Our methodology can easily be used for assessing such models.
@@
Another extension to this methodology is to take more steps along the gradient flow for a given objective. It may be feasible to learn such a transformation using neural or other expressive representations. Finally, we resolved the spectrum of the Hessian matrix for shallow ReLU models – the spectral decomposition for larger ReLU models may be useful for other analyses.
Acknowledgements